OpenStudy (anonymous):

Write the standard equation of the circle with center (-8, -7) and a radius of 6.

OpenStudy (anonymous):

A) (x + 8)2 + (y + 7)2 = 6 B) (x – 8)2 + (y – 7)2 = 6 C) (x + 8)2 + (y + 7)2 = 36 D) (x – 8)2 + (y – 7)2 = 36

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

if centre is \((a,b)\) and radius is \(r\) the equation of the circle is \((x-a)^2+(y-b)^2=r^2\)

OpenStudy (anonymous):

but there is no equation that adds up to that!!!

OpenStudy (triciaal):

yes there is

OpenStudy (anonymous):

No it is either both subtraction or both addition

OpenStudy (triciaal):

when you subtract a negative it becomes a positive

OpenStudy (triciaal):

now try again

OpenStudy (anonymous):

So it would be B?

OpenStudy (triciaal):

no

OpenStudy (anonymous):

It would be A? Because the equation has to add up to 6

OpenStudy (triciaal):